# Trust or Bust: Ensuring Trustworthiness in Autonomous Weapon Systems

### v1

@@ Abstract
The integration of Autonomous Weapon Systems (AWS) into military operations presents both significant opportunities and challenges. This paper explores the multifaceted nature of trust in AWS, emphasising the necessity of establishing reliable and transparent systems to mitigate risks associated with bias, operational failures, and accountability. Despite advancements in Artificial Intelligence (AI), the trustworthiness of these systems, especially in high-stakes military applications, remains a critical issue. Through a systematic review of existing literature, this research identifies gaps in the understanding of trust dynamics during the development and deployment phases of AWS. It advocates for a collaborative approach that includes technologists, ethicists, and military strategists to address these ongoing challenges. The findings underscore the importance of Human-Machine teaming and enhancing system intelligibility to ensure accountability and adherence to International Humanitarian Law. Ultimately, this paper aims to contribute to the ongoing discourse on the ethical implications of AWS and the imperative for trustworthy AI in defense contexts.
@@ II Related Work
In the context of AI, autonomy refers to a system’s ability to make decisions and act independently without human intervention. This capability extends beyond simple automation, as autonomous AI systems can adapt their actions based on learning from their environment and situational awareness. For AWS, autonomy is characterised by the ability to select and engage targets independently, based on predefined criteria and environmental interpretation. AWS operate in complex military environments requiring rapid decision-making. They use self-learning capabilities to adapt over time, navigating unpredictable scenarios without constant reprogramming. Nevertheless, the definition of AWS also considers the degree of human control, which varies among different interpretations. Some emphasise maintaining human oversight in critical decision-making processes, while others focus on the system’s independent operation. This balance between autonomy and human control is a central point of debate in the ethical and legal considerations surrounding AWS [7]. To address this, [8] proposes adopting a systemic view when building and using AWS. This perspective involves a comprehensive approach that recognises AWS as complex entities within larger systems supported by three fundamental pillars: system, autonomy, and values. The system pillar views AWS as self-regulating components interacting with human operators, legal frameworks, and ethical considerations. The autonomy pillar highlights AWS’s independent operation, making decisions based on programming and data. The values pillar addresses the legal and moral considerations governing AWS use, ensuring compliance with International Humanitarian Law (IHL) and ethical standards. This holistic approach acknowledges the inherent complexity of AWS and the challenges in their regulation. By adopting a systemic vision, stakeholders can better understand the technological, operational, legal, and ethical dimensions of AWS, leading to more effective development, deployment, and regulation strategies.
@@
The regulation of AWS is an evolving field that addresses the rapid advancements in military technology and their ethical, legal, and security implications. Key aspects of regulation include the need for updated and comprehensive guidelines that go beyond existing frameworks. The Advisory Council on International Affairs and the Advisory Committee on Issues of Public International Law advocate for more concrete regulations that cover the development, procurement, and use of AWS, emphasising the importance of clear criteria and responsibilities for all stakeholders involved. Additionally, international cooperation involving intensified consultations among governments, the private sector, and civil society is necessary to ensure that AWS development remains within ethical and legal boundaries. The involvement of a broad range of stakeholders is crucial to creating effective, widely accepted, regulatory frameworks that address the potential for abuse and ensure compliance with IHL [9, 10, 11].
@@
In particular, Lethal Autonomous Weapon Systems (LAWS) raise several core issues that are central to ongoing debates. Ethically, the primary concern is the moral acceptability of machines making life-and-death decisions, which challenges the principles of distinction and proportionality required by the law of war. Accountability is another major issue, as determining responsibility for unlawful actions or malfunctions becomes complex when autonomous systems are involved. Operational risks, including the potential for malfunctions or exploitation by adversaries, further complicate the deployment of LAWS, necessitating robust safeguards to minimise failures. Compliance with international law is also critical, as there are significant doubts about whether AWS can adhere to legal standards in complex combat scenarios. Finally, the absence of international consensus on LAWS’ definition and regulation hinders control measures, given diverse views on their usage and the possibility of a ban or strict rules. These issues reveal the need for careful consideration and comprehensive regulation to address the ethical, legal, and operational challenges posed by LAWS [4, 1].
@@
A perspective that allows tackling some of the existing issues is forming and engaging in human-autonomy teaming (HAT). In the context of military operations, this refers to the collaboration between human operators and autonomous systems to achieve common goals, leveraging the strengths of both. This concept is crucial for enhancing operational capabilities, as autonomous systems can process large volumes of data rapidly and operate in hazardous environments, providing superior situational awareness and decision-making support. Trust between humans and autonomous systems is foundational to effective HAT, ensuring that operators are comfortable relying on AI for proactive decisions in dynamic battlefield conditions. Cultural and psychological factors, such as automation bias, influence trust and must be understood to maximise the benefits of AI integration. Additionally, HAT must operate within strategic, political, and legal frameworks, adhering to international laws of armed conflict. The continuous advancement of AI technology offers new opportunities for improving military operations, making the successful implementation of HAT essential for maintaining operational effectiveness and strategic advantage [12, 13].
@@
It is then important, as [14, 15] argue, to understand both the challenges and opportunities as well as the disadvantages and advantages of building and using AWS. On the positive side, AWS drive innovation, enhance defence capabilities, and offer strategic advantages by operating in high-risk environments with increased precision and speed. They have the potential to reduce human casualties and increase operational efficiency through rapid data processing and independent task execution. However, AWS also face significant challenges, including technological and financial barriers, cyber security risks, and critical moral and legal concerns. The ethical implications of entrusting life-and-death decisions to machines, potential violations of international laws of war, and issues of accountability are paramount. Additionally, the rapid pace of technological advancement complicates effective regulation and definition of AWS, while potentially exacerbating inequalities between nations due to dependencies on foreign technology. Balancing these pros and cons is essential for responsible development and deployment, requiring careful consideration of ethical, legal, and operational implications to ensure alignment with international laws and moral standards. Hence, in current as well as in future applications, the proper governance of AWS is crucial to address responsibly and safely the complex interplay of ethical, legal, security, and technological challenges they present, both now and in future applications. Effective governance frameworks are essential to ensure that AWS operate within established ethical and legal boundaries, particularly concerning accountability and compliance with IHL. As AWS have the potential to transform military strategies, governance is necessary to manage the risks associated with their deployment, including escalation of conflicts and the proliferation of these technologies to malicious actors. Moreover, the governance of AWS must account for the political and economic implications of their development, balancing national security interests with the potential for destabilisation in international relations. Establishing clear guidelines for the production, deployment, and use of AWS is vital to maintain human oversight and accountability, ensuring that decisions regarding lethal force remain under human control [16, 17].
@@
Nevertheless, as this extensive literature review shows, while various academic and practitioner studies are dedicated to capturing various dimensions characterising the processes of building and using AWS in military operations, a systematic approach to understand and be able to further materialise corresponding challenges and opportunities in the military domain is lacking. This represents the knowledge gap that this article aims to tackle in this study.
@@ IV Challenges of AWS
Challenge 2 - Training and Adaptability: Another key challenge is managing the balance of trust between humans and AWS. It is crucial to ensure that human actors neither over- nor under-trust these systems. Despite their exceptional adaptability, humans can often misjudge a machine’s capabilities, leading to ‘automation bias’. This bias can manifest in two ways: humans might over-trust the system, accepting its decisions even when they are unjust, or they might under-trust it, constantly monitoring or overriding the system’s tasks  [19, 20, 31].
@@
Challenge 3 - Strategic vulnerability: One of the key issues is the concept of first-strike instability, which emerges as a strategic vulnerability in the development and deployment of AWS. This instability arises from the inherent characteristics of AWS, such as their speed, precision, and the ability to limit human risk. These attributes potentially incentivize a preemptive strike, offering tactical advantages like rapid and accurate attacks at a lower cost, without endangering personnel. This scenario, where one party is motivated to strike first due to these advantages, constitutes first-strike instability, even as AI continues to evolve towards reduced brittleness and increased transparency  [63]. This issue of first-strike instability is further complicated by the secretive nature of AWS, which can lead to increased uncertainty among allied nations. Specifically, nations might be reluctant to share details about their AWS capabilities, which could potentially put a strain on their relationships  [63]. On a practical level, there are doubts about the ability of AI-enabled systems to accurately differentiate between legitimate and illegitimate targets. Moreover, the potential consequences of any misidentification are immense, further amplifying the complexity of the issue at hand [40, 26]. The first-strike instability of AWS presents several challenges. Decision-making algorithms often lack transparency, leaving policymakers in the dark—a problem known as the ‘black box’ challenge. Additionally, AWS faces data-related issues such as data acquisition, data skew, misuse, and privacy concerns. At the same time, the potential of AI to perpetuate existing discriminatory structures and biases is a significant concern that needs to be addressed. Lastly, the role of humans in these autonomous systems varies across contexts; in military systems, human oversight may be legally or ethically required, whereas, in transportation, human involvement might be less necessary - especially if human error is a major contributor to operational incidents [39]. In the context of strategic vulnerability, accountability and responsibility are major challenges. It is essential to establish clear rules and mechanisms when using AWS to prevent irresponsible behaviour. However, attempts to translate laws into software have proven to be challenging, and no practical strategy exists to evaluate proportionality in complex military scenarios [54].
@@
Challenge 5 - Capturing ethical, legal, and social dimensions: The push for incorporating ethical principles in AWS highlights the unresolved task of choosing the most fitting ethical standard for a situation. Various ethical theories have been used to tackle moral dilemmas. However, there is currently no consensus among ethical experts regarding the choice of ethical theory to apply [61, 46, 26]. Establishing robust accountability mechanisms is essential when deploying AWS to prevent systemic irresponsibility. The risk of AI amplifying discriminatory structures is a significant concern, necessitating proactive measures to counteract potential bias [54, 39]. Furthermore, the societal implications and public perception of AWS also present a significant challenge. Central issues include the AI potential to reinforce existing discriminatory structures and the democratic accountability of autonomous systems, particularly when these systems make decisions affecting public rights or the allocation of public benefits [54, 44].
Subsequently, transparency and explainability are also significant challenges. A primary concern is the transparency of autonomous systems in terms of decision-making, often referred to as the ‘black box’ challenge. Unlike the human brain, which continuously learns and builds upon previously acquired knowledge, algorithms are typically trained once on a dataset. This approach restricts their ability to learn new information without retraining, posing challenges in dynamic and complex environments where continuous learning is vital  [41, 31].
Finally, using autonomous systems in military contexts brings ethical dilemmas related to human rights and warfare principles into focus. The ability of humans to challenge the codified ethics of these machines could be reduced, leading to a moral vacuum where existing laws, ethical guidelines, and societal norms may be insufficient. On a more pragmatic note, there are doubts about whether AI-driven functionalities can accurately distinguish between lawful and unlawful targets  [41, 43].
@@ V Opportunities of AWS
Opportunity 3 - Strategic and Military Advantage: The development and integration of AWS into military operations present significant strategic advantages for technologically advanced nations, particularly in enhancing precision, speed, and efficiency [19, 20]. This focus on machine autonomy is driven by the potential for improved operational effectiveness through Human-Machine teaming [20]. AWS integration into C2 systems offers the prospect of enhanced decision-making processes, providing commanders with data-driven insights and a broader range of tactical options [39, 42]. Public perception of AWS appears to be malleable, with increased support when presented with information highlighting their potential to save soldiers’ lives and reduce errors [39]. AWS demonstrate enhanced combat efficiency and risk mitigation by operating in high-threat environments unsuitable for human soldiers, thus increasing operational capabilities while reducing risk to human life [39]. These systems act as force multipliers, performing tasks that would otherwise require multiple human soldiers, thereby improving operational efficiency [59, 62]. European weapon manufacturers have contributed to cumulative improvements in AWS through design choices and practices, shaping implicit rules for appropriate human control in specific use-of-force situations [48, 49, 50].

### v2

@@ Abstract
The integration of Autonomous Weapon Systems (AWS) into military operations presents both significant opportu- nities and challenges. This paper explores the multifaceted nature of trust in AWS, emphasising the necessity of establishing reliable and transparent systems to mitigate risks associated with bias, operational failures, and accountability. Despite advancements in Artificial Intelligence (AI), the trustworthiness of these systems, especially in high-stakes military applications, remains a critical issue. Through a systematic literature review, this research identifies gaps in the understanding of trust dynamics during the development and deployment phases of AWS. It advocates for a collaborative approach that includes technologists, ethicists, and military strategists to address these ongoing challenges. The findings underscore the importance of Human-Machine teaming and enhancing system intelligibility to ensure accountability and adherence to law. Ultimately, this paper aims to contribute to the ongoing discourse on the ethical implications of AWS and the imperative for trustworthy AI in defense contexts.
@@ II Related Work
In the context of AI, autonomy refers to a system’s ability to make decisions and act independently without human intervention. This capability extends beyond simples automation, as autonomous AI systems can adapt their actions based on learning from their environment and situational awareness. For AWS, autonomy is characterised by the ability to select and engage targets independently, based on predefined criteria and environmental interpretation. AWS operate in complex environments requiring rapid decision-making. They use self-learning capabilities to adapt over time, navigating unpredictable scenarios without constant reprogramming. Nevertheless, the definition of AWS also considers the degree of human control, which varies among different interpretations. Some emphasise maintaining human oversight in critical decision-making processes, while others focus on the system’s independent operation. Balance between autonomy and human control is a central point of debate in the ethical and legal considerations surrounding AWS [7].
@@
To address this, [8] proposes adopting a systemic view when building and using AWS. This involves a comprehensive approach that recognises AWS as complex entities within larger systems supported by three fundamental pillars: system, autonomy, and values. The system pillar views AWS as self-regulating components interacting with humans, laws and ethics. The autonomy pillar highlights AWS’s independent operation based on programming and data. The values pillar addresses the legal and moral considerations governing AWS use, ensuring compliance with International Humanitarian Law (IHL) and ethical standards. This holistic approach acknowledges the inherent complexity of AWS and the challenges in their regulation. By adopting a systemic vision, stakeholders can better understand the technological, operational, legal, and ethical dimensions of AWS, leading to more effective development, deployment, and regulation strategies. Key aspects of regulation include the need for updated and comprehensive guidelines that go beyond existing frameworks. The Advisory Council on International Affairs and the Advisory Committee on Issues of Public International Law advocate for more concrete regulations that cover the development, procurement, and use of AWS, emphasising the importance of clear criteria and responsibilities for all stakeholders involved. Additionally, international cooperation involving intensified consultations among governments, the private sector, and civil society is necessary to ensure that AWS development remains within ethical and legal boundaries. The involvement of a broad range of stakeholders is crucial to creating effective, widely accepted, regulatory frameworks that address the potential for abuse and ensure compliance with IHL [9, 10, 11].
@@
In particular, Lethal Autonomous Weapon Systems (LAWS) pose ethical, legal, and operational challenges. Ethically, the primary concern is the moral acceptability of machines making life-and-death decisions, which challenges the principles of distinction and proportionality required by the law of war. Accountability is another major issue, as determining responsibility for unlawful actions or malfunctions becomes complex when autonomous systems are involved. Operational risks, including the potential for malfunctions or exploitation by adversaries, further complicate the deployment of LAWS, necessitating robust safeguards to minimise failures. Compliance with international law is also critical, as there are significant doubts about whether AWS can adhere to legal standards in complex combat scenarios. Finally, the absence of international consensus on LAWS’ definition and regulation hinders control measures, given diverse views on their usage and the possibility of a ban or strict rules. These issues reveal the need for careful consideration and comprehensive regulation to address the ethical, legal, and operational challenges posed by LAWS [4, 1].
@@
A perspective that allows tackling some of the existing issues is forming and engaging in human-autonomy teaming (HAT). In the context of military operations, this refers to the collaboration between human operators and autonomous systems to achieve common goals, leveraging the strengths of both. This concept is crucial in enhancing operational capabilities, as autonomous systems can process large volumes of data rapidly and operate in hazardous environments, providing superior situational awareness and decision-making support. Trust between humans and autonomous systems is foundational to effective HAT, ensuring that operators are comfortable relying on AI for proactive decisions in dynamic battlefield conditions. Cultural and psychological factors, such as automation bias, influence trust and must be understood to maximise the benefits of AI integration. Additionally, HAT must operate within strategic, political, and legal frameworks, adhering to international laws of armed conflict. The continuous advancement of AI offers new opportunities, making the successful implementation of HAT essential for maintaining operational effectiveness and strategic advantage [12, 13].
@@
It is then important, as [14, 15] argue, to understand both the challenges and opportunities as well as the (dis)advantages of building and using AWS. On the positive side, AWS drive innovation, enhance defence capabilities, and offer strategic advantages by operating in high-risk environments with increased precision and speed. They have the potential to reduce human casualties and increase operational efficiency through rapid data processing and independent task execution. However, AWS also face significant challenges, including technological and financial barriers, cyber security risks, and critical moral and legal concerns. The ethical implications of entrusting life-and-death decisions to machines, potential violations of international laws of war, and issues of accountability are paramount. Additionally, the rapid pace of technological advancement complicates effective regulation and definition of AWS, while potentially exacerbating inequalities between nations due to dependencies on foreign technology. Balancing these pros and cons is essential for responsible development and deployment, requiring careful consideration of ethical, legal, and operational implications to ensure alignment with international laws and moral standards. Hence, in current and future applications, proper governance of AWS is crucial to address responsibly and safely the complex interplay of ethical, legal, security, and technological challenges they present, both now and in future applications. Effective governance frameworks are crucial to ensure AWS operate ethically and legally, especially in terms of accountability and compliance with IHL. As AWS has the potential to transform military strategies, governance is needed to manage the risks associated with their deployment, including escalation of conflicts and the proliferation of these technologies to malicious actors. Moreover, governance of AWS must account for the political and economic implications of their development, balancing national security interests with the potential for destabilisation in international relations. Establishing clear guidelines for the production, deployment, and use of AWS is vital to maintain human oversight and accountability, ensuring decisions regarding lethal force remain under human control [16, 17].
@@
Nevertheless, as this extensive review shows, while various academic and practitioner studies are dedicated to capturing various dimensions characterising the processes of building and using AWS in military operations, a systematic approach to understand and be able to further materialise corresponding challenges and opportunities in the military domain is lacking. This represents the knowledge gap that this article aims to tackle in this study.
@@ IV Challenges of AWS
Challenge 2 - Training and Adaptability: Another key challenge is managing the balance of trust between humans and AWS. It is crucial to ensure that human actors neither over- nor under-trust these systems. Despite their exceptional adaptability, humans often misjudge a machine’s capabilities, leading to ‘automation bias’. This can manifest in two ways: humans might over-trust the system, accepting its decisions even when they are unjust, or they might under-trust it, constantly monitoring or overriding the system’s tasks  [19, 20, 31].
@@
Challenge 3 - Strategic vulnerability: One of the key issues is the concept of first-strike instability, which emerges as a strategic vulnerability in the development and deployment of AWS. This instability arises from the inherent characteristics of AWS, such as their speed, precision, and the ability to limit human risk. These attributes potentially incentivize a preemptive strike, offering tactical advantages like rapid and accurate attacks at a lower cost, without endangering personnel. This scenario, where one party is motivated to strike first due to these advantages, constitutes first-strike instability, even as AI continues to evolve towards reduced brittleness and increased transparency  [63]. This issue of first-strike instability is further complicated by the secretive nature of AWS, which can lead to increased uncertainty among allied nations. Specifically, nations might be reluctant to share details about AWS capabilities, which could potentially put a strain on relationships  [63]. On a practical level, there are doubts about the ability of AI-enabled systems to accurately differentiate between legitimate and illegitimate targets. Moreover, the potential consequences of any misidentification are immense, further amplifying the complexity of the issue at hand [40, 26]. The first-strike instability of AWS presents several challenges. Decision-making algorithms often lack transparency, leaving policymakers in the dark - a problem known as the ‘black box’ challenge. Additionally, AWS faces data-related issues such as data acquisition, data skew, misuse, and privacy concerns. At the same time, the potential of AI to perpetuate existing discriminatory structures and biases is a significant concern that needs to be addressed. Lastly, the role of humans in these autonomous systems varies across contexts; in military systems, human oversight may be legally or ethically required, whereas, in transportation, human involvement might be less necessary - especially if human error is a major contributor to operational incidents [39]. In the context of strategic vulnerability, accountability and responsibility are major challenges. It is essential to establish clear rules and mechanisms when using AWS to prevent irresponsible behaviour. However, attempts to translate laws into software have proven to be challenging, and no practical strategy exists to evaluate proportionality in complex military scenarios [54].
@@
Challenge 5 - Capturing ethical, legal, and social dimensions: The push for incorporating ethical principles in AWS highlights the unresolved task of choosing the most fitting ethical standard for a situation. Various ethical theories have been used to tackle moral dilemmas. However, there is currently no consensus among ethical experts regarding the choice of ethical theory to apply [61, 46, 26]. Establishing robust accountability mechanisms is essential when deploying AWS to prevent systemic irresponsibility. The risk of AI amplifying discriminatory structures is a significant concern, necessitating proactive measures to counteract potential bias [54, 39]. Furthermore, the societal implications and public perception of AWS also present a significant challenge. Central issues include the AI potential to reinforce existing discriminatory structures and the democratic accountability of autonomous systems, particularly when these systems make decisions affecting public rights or the allocation of public benefits [54, 44]. Subsequently, transparency and explainability are also significant challenges. A primary concern is the transparency of autonomous systems in terms of decision-making, often referred to as the ‘black box’ challenge. Unlike the human brain, which continuously learns and builds upon previously acquired knowledge, algorithms are typically trained once on a dataset. This approach restricts their ability to learn new information without retraining, posing challenges in dynamic and complex environments where continuous learning is vital [41, 31]. Finally, using autonomous systems in military contexts brings ethical dilemmas related to human rights and warfare principles into focus. The ability of humans to challenge the codified ethics of these machines could be reduced, leading to a moral vacuum where existing laws, ethical guidelines, and societal norms may be insufficient. On a more pragmatic note, there are doubts about whether AI-driven functionalities can accurately distinguish between lawful and unlawful targets  [41, 43].
@@ V Opportunities of AWS
Opportunity 3 - Strategic and Military Advantage: The development and integration of AWS into military operations present significant strategic advantages for technologically advanced nations, particularly in enhancing precision, speed, and efficiency [19, 20]. This focus on machine autonomy is driven by the potential for improved operational effectiveness through Human-Machine teaming [20]. AWS integration into C2 systems offers the prospect of enhanced decision-making processes, providing commanders with data-driven insights and a broader range of tactical options [39, 42]. Public perception of AWS appears to be malleable, with increased support when presented with information highlighting their potential to save soldiers’ lives and reduce errors [39].AWS enhance combat efficiency and mitigate risk by operating in high-threat environments, boosting capabilities while reducing risk to human life [39]. These systems act as force multipliers, performing tasks that would otherwise require multiple human soldiers, thereby improving operational efficiency [59, 62]. European weapon manufacturers have contributed to cumulative improvements in AWS through design choices and practices, shaping implicit rules for appropriate human control in specific use-of-force situations [48, 49, 50].